Evaluate: 3 sov 16 de (4-22)5/2 B IV AA - Ix E 2 3 x TTT Р
Answer #1

# substitutions - 16 da (4-x2) 5/2 0 x=2 sin o da= 2 Cosodo. 16 dx 2 los o do = sino (A - 4 sino] / o= sint (2) 0 n 32 Coso d